# The Boundaries of Tractability in Hierarchical Task Network Planning

### v1

@@ 2 Preliminaries
Input: A task network $tn$ over some domain $(F,A,{\cal C},\delta,M)$ and an initial state $s_{0}\subseteq F$.
Question: Is there an executable plan starting in $s_{0}$ that executes all the tasks in $tn$?
@@
Input: A task network $tn$ over some domain $(F,A,{\cal C},\delta,M)$, an initial state $s_{0}\subseteq F$, and a multiset $S$ of actions in $A$.
Question: Is there a plan (not necessarily using all available tasks) that is executable in $tn$ from $s_{0}$ and that covers $S$?
Action Executability is a generalization of Plan Existence in the sense that, on primitive task networks, Plan Existence simply equals Action Executability if $S$ is the multiset of actions obtained by mapping each task by $\alpha$. On general networks, any Plan Existence instance can be reduced to Action Executability by adding a new task $t$ with new action $a$, letting $S={\{a\}}$, and adding $(t^{\prime},t)$ to $\prec$ for each sink $t^{\prime}$ in the task network. As this may change some structural parameters associated to the task network, alternatively one can create a new task $t_{t^{\prime}}$ with action $a$ for each sink $t^{\prime}$, only add the edges $(t^{\prime},t_{t^{\prime}})$ to $\prec$, and let $S$ contain $a$ as many times as there are sinks. Therefore, algorithmic and hardness results transfer in the respective directions; in our setting, we establish all of our hardness results for the “simpler” Plan Existence, while all of our algorithms can deal with the “harder” Action Executability.
@@
Input: A task network $tn$ over some domain $(F,A,{\cal C},\delta,M)$, an initial state $s_{0}\subseteq F$, and a target state $s_{g}\subseteq F$.
Question: Is there a plan (not necessarily using all available tasks) that is executable in $tn$ from $s_{0}$ and that results in a state $s\supseteq s_{g}$?
We remark that, while Plan Validation, Plan Existence, Action Executability, and State Reachability are stated as decision problems for complexity-theoretic purposes, all of our obtained algorithms are constructive and can output a task sequence or plan as a witness. For each of these problems, we define the input size to simply be the sum of the sizes of elements occurring on the input; for example, for an instance ${\cal I}$ of Plan Validation this would be $|{\cal I}|=|T|+|\prec^{+}|+|\alpha|+|F|+|A|+|{\cal C}|+|\delta|+|M|$.
For some of our algorithms, we employ a state transition graph. For a set of actions $A$ and an initial state $s_{0}$, we define the state transition graph to be a multigraph ${\cal G}=({\cal S},{\cal T},\beta)$ that describes the set ${\cal S}$ of states reachable from $s_{0}$ and how actions transform one state into another. We construct ${\cal G}$ inductively as follows. First, initialize ${\cal S}={\{s_{0}\}}$ and ${\cal T}=\emptyset$. Then, update from $s_{0}$. To update from a state $s$, iterate through all actions $a\in A$. If $\operatorname{prec}(a)\subseteq s$, let $s^{\prime}=(s\setminus\operatorname{del}(a))\cup\operatorname{add}(a)$. If $s^{\prime}\notin{\cal S}$, add $s^{\prime}$ to ${\cal S}$ and update from $s^{\prime}$. In any case, add $e=(a,s,s^{\prime})$ to ${\cal T}$ and let $\beta(e)=(s,s^{\prime})$. The computation is completed once no more updates occur. An example is provided in Fig. 3. This construction takes at most $\mathcal{O}\!\left(|{\cal I}|^{2}\right)$ time. Observe that ${\cal G}$ contains exactly all states that $s_{0}$ can be transformed into by applying sequences of actions in $A$. Further, a pair of states is connected by directed edges representing all actions that transform one of the states into the other. It is known that the number of states $|{\cal S}|$ can be upper-bounded by $|A|!\cdot 2^{|A|}$ Kronegger et al. (2013).
$a_{1}$emptyset $\emptyset$ ${\{1\}}$
$a_{2}$${\{1\}}$emptyset ${\{2\}}$
$a_{3}$${\{2\}}$${\{2,3\}}$emptyset
$a_{4}$${\{1\}}$emptyset ${\{2,3\}}$
$a_{5}$${\{2\}}$emptyset ${\{3\}}$
Figure 3: Example of a state transition graph with initial state $s_{0}=\emptyset$ and actions $A={\{a_{1},a_{2},a_{3},a_{4},a_{5}\}}$. An edge labeled with $a_{i}$ that connects some state $s$ to some state $s^{\prime}$ refers to the edge $(a_{i},s,s^{\prime})$ in the state transition graph. An edge labeled with multiple actions refers to multiple edges, one for each action, in the multigraph.
Parameterized Complexity.
Parameterized complexity Downey and Fellows (2013); Cygan et al. (2015) analyzes the running time of algorithms with respect to a parameter $k\in\mathbb{N}$ and input size $n$. The high-level idea is to find a parameter that describes the structure of the instance such that the combinatorial explosion can be confined to this parameter. In this respect, the most favorable complexity class is FPT (fixed-parameter tractable) which contains all problems that can be decided by an algorithm running in $f(k)\cdot n^{\mathcal{O}(1)}$ time, where $f$ is a computable function. Algorithms with this running-time are called fixed-parameter algorithms.
Showing that a problem is hard for the complexity classes ${\textsf{W}}[1]$ or ${\textsf{W}}[2]$ rules out the existence of a fixed-parameter algorithm under the well-established assumption that ${\textsf{W}}[1]\neq{\textsf{FPT}}$. This is done via a parameterized reduction Downey and Fellows (2013); Cygan et al. (2015) from some known ${\textsf{W}}[1]$-hard (${\textsf{W}}[2]$-hard, respectively) problem. A parameterized reduction from a parameterized problem $\mathcal{P}$ to a parameterized problem $\mathcal{Q}$ is a function:
  • which maps yes-instances to yes-instances and no-instances to no-instances,
  • which can be computed in $f(k)\cdot n^{\mathcal{O}(1)}$ time, where $f$ is a computable function, and
  • where the parameter of the output instance can be upper-bounded by some function of the parameter of the input instance.
3 Setting the Stage: The State Space
As a basic precondition for solving our problems on more advanced task networks, we need to ensure tractability at least on “trivial” primitive task networks where $\prec^{+}=\emptyset$. Unfortunately, in their most general form, all but the first of our problems of interest turn out to be intractable even on these degenerate networks via a simple reduction from 3-SAT.
@@ 3 Setting the Stage: The State Space
State Reachability and Plan Existence are NP-hard on primitive task networks with $\prec^{+}\ =\emptyset$.
@@
We give a reduction from 3-SAT to State Reachability. Given an instance $\phi$ of 3-SAT, we create an instance $(tn,s_{0},s_{g})$ of State Reachability as follows. For each clause $C_{j}$ in $\phi$, $j\in[m]$, we create a proposition $C_{j}$ in $F$. For each variable $x_{i}$ in $\phi$, $i\in[n]$, we create a proposition $x_{i}$ in $F$, as well as two actions $a_{x_{i}}$, $a_{\overline{x}_{i}}$ and two tasks $t_{x_{i}}$, $t_{\overline{x}_{i}}$ that are mapped to $a_{x_{i}}$ and $a_{\overline{x}_{i}}$, respectively. For all $i\in[n]$, let $C_{x_{i}}$ ($C_{\overline{x}_{i}}$, respectively) be the set of all propositions of the form $C_{j}$ that correspond to clauses in $\phi$ that are satisfied by setting the variable $x_{i}$ to TRUE (FALSE, respectively). For all $i\in[n]$, we set $\operatorname{prec}(a_{x_{i}})=\operatorname{prec}(a_{\overline{x}_{i}})=\operatorname{del}(a_{x_{i}})=\operatorname{del}(a_{\overline{x}_{i}})=\{x_{i}\}$, $\operatorname{add}(a_{x_{i}})=C_{x_{i}}$, and $\operatorname{add}(a_{\overline{x}_{i}})=C_{\overline{x}_{i}}$. Finally, we set $s_{0}=\{x_{1},\ldots,x_{n}\}$ and $s_{g}=\{C_{1},\ldots,C_{m}\}$. Note that $tn$ is primitive with $\prec^{+}=\emptyset$. The equivalence between the two instances is easy to see. Indeed, since $s_{0}=\{x_{1},\ldots,x_{n}\}$ and $\operatorname{prec}(a_{x_{i}})=\operatorname{prec}(a_{\overline{x}_{i}})=\operatorname{del}(a_{x_{i}})=\operatorname{del}(a_{\overline{x}_{i}})=\{x_{i}\}$, only one of the two tasks $t_{x_{i}}$, $t_{\overline{x}_{i}}$ may be executed for each $i\in[n]$. Further, the execution of $t_{x_{i}}$ ($t_{\overline{x}_{i}}$, respectively) causes the execution of $a_{x_{i}}$ ($a_{\overline{x}_{i}}$, respectively), which adds, to the current state, all the propositions corresponding to clauses in $\phi$ that are satisfied by setting the variable $x_{i}$ to TRUE (FALSE, respectively). Finally, $s_{g}$ contains all the clause propositions. Hence, $\phi$ is satisfiable if and only if $(tn,s_{0},s_{g})$ is a yes-instance of State Reachability. We can also show that the same holds for Action Executability, even if $|S|=1$; indeed, it suffices to use the same construction with the addition of a new task $t_{a^{*}}$ that is mapped to a new action $a^{*}$ with $S=\{a^{*}\}$ and $\operatorname{prec}(a^{*})=\{C_{1},\ldots,C_{m}\}$. Finally, for Plan Existence, we begin by recalling that the problem is NP-hard on primitive task networks (Erol et al., 1996, Theorem 8). We can reduce from Plan Existence on primitive task networks to Plan Existence on primitive task networks with $\prec^{+}=\emptyset$ by encoding the partial order into the preconditions. Let there be a task network $tn=(T,\prec^{+},\alpha)$ with action set $A$ and set of propositions $F$. Then, construct a new task network $tn^{\prime}=(T,\emptyset,\alpha^{\prime})$ with action set $A^{\prime}$ and proposition set $F^{\prime}$ as follows. Let $F^{\prime}=F\cup{\{f_{t}\mid t\in T\}}$. For each $t\in T$, create an action $a_{t}\in A^{\prime}$ such that $\operatorname{prec}(a_{t})=\operatorname{prec}(\alpha(t))\cup{\{f_{t^{\prime}}\mid(t^{\prime},t)\in\prec^{+}\}}$, $\operatorname{del}(a_{t})=\operatorname{del}(\alpha(t))$, and $\operatorname{add}(a_{t})=\operatorname{add}(\alpha(t))\cup{\{f_{t}\}}$. Let $\alpha^{\prime}(t)=a_{t}$ for all $t\in T$. Observe that this way, a task in $tn^{\prime}$ can only be executed if the respective preconditions as defined for the original instance are fulfilled and all its predecessors in $\prec^{+}$ are executed. Thus, any task sequence is executable in $tn^{\prime}$ from some initial state $s_{0}$ if and only if it is executable in $tn$ from $s_{0}$. ∎
A crucial requirement for the reductions underlying Theorem 1 is that the state space (i.e., the number of states) can be large, and in particular not bounded by any constant. As we will see in the next sections, tractability can be achieved if the state space is assumed to be fixed (that is, bounded by a constant that is independent of the input). Hence, we will hereinafter analyze the affected problems (State Reachability, Plan Existence, and its generalization Action Executability) in the bounded state-space setting.
4 Primitive Task Networks
Our first set of results focuses on the classical complexity of the targeted problems on primitive task networks. We begin with a series of lower bounds which rule out polynomial-time algorithms even for some very simple “tree-like” networks.
The Intractability of Tree-like HTNs.
First, note that the intractability of Plan Validation on tree-like networks was already established by previous work of Lin and Bercher:
Theorem 2 (Lin and Bercher (2023), Theorem 3).
Plan Validation on primitive task networks is NP-hard even if $G_{\prec}$ is a forest of stars and $\operatorname{prec}(a)=\operatorname{del}(a)=\operatorname{add}(a)=\emptyset$ for all actions $a\in A$.
Further, Plan Validation remains intractable even if $\prec^{+}$ describes a set of total orders that are independent from one another (i.e., if $D_{\prec}$ is a disjoint union of paths). We establish this via a reduction from the Shuffle Product problem:
Shuffle Product
Input: Words $u,c_{1},\ldots,c_{w}$ over an alphabet $\Sigma$.
Question: Can $u$ be generated by interlacing the words in $c_{1},\ldots,c_{w}$, i.e., can the letters of $c_{1},\ldots,c_{w}$ be rearranged to form $u$ while preserving the order the letters appear in for each individual word $c_{i}$, $i\in[w]$?
Shuffle Product is known to be NP-hard even when restricted to the case where $|\Sigma|=2$ Warmuth and Haussler (1984).
@@ 4 Primitive Task Networks
Plan Validation is NP-hard, even when restricted to primitive task networks such that $\prec^{+}$ is a collection of chains, $|A|=2$, and $\operatorname{prec}(a)=\operatorname{del}(a)=\operatorname{add}(a)=\emptyset$ for all actions $a\in A$.
@@
We reduce from Shuffle Product with the alphabet $\Sigma={\{a,b\}}$. Let $A=\Sigma$. Construct a task for each letter in the words $c_{1},\ldots,c_{w}$. In particular, let the task $t_{i,j}$ refer to the $j^{\text{th}}$ letter of the word $c_{i}$, $i\in[w]$. Let $\prec^{+}$ be such that the tasks for each word form a chain in the respective order, i.e., let $\prec$ consist of the edges $(t_{i,j},t_{i,j+1})$, for all $i\in[w]$ and $j\in[|c_{i}|-1]$. Then, $\prec^{+}$ partitions $T$ into $w$ chains (see Fig. 4). Now, set $\alpha(t_{i,j})$ to the action representing the $j^{\text{th}}$ letter of $c_{i}$. Let the input plan $p$ be such that $p=u$. The constructed instance consists of $w$ chains and the construction takes polynomial time.
The correctness of the reduction follows from the fact that verifying whether the tasks in this Plan Validation instance can be arranged to match the input plan corresponds to interlacing the words to form $u$. ∎
A different reduction from Shuffle Product yields similar hardness results for the other problems; it will be useful to recall that hardness for Plan Existence directly carries over to hardness for Action Executability.
@@
State Reachability and Plan Existence are NP-hard, even when restricted to primitive task networks such that $\prec^{+}$ is a collection of chains, $|A|\leq 5$, and the state transition graph has at most $4$ states.

### v2

@@ 2 Preliminaries
Input: A task network $tn$ in a domain $D$ and an initial state $s_{0}\subseteq F$.
Question: Is there a full solution to $(D,tn,s_{0})$?
In addition to the problems mentioned in the introduction, we also define the corresponding problem of constructing a plan executing (at least) a specified set of actions. This Action Executability problem is closely related to Plan Existence and, in fact, all results that we prove for one of the problems will also hold for the other.
@@
Input: A task network $tn$ in a domain $D$, an initial state $s_{0}\subseteq F$, and a multiset $S$ of actions in $A$.
Question: Is there a solution to $(D,tn,s_{0})$ with a linearization $\bar{t}$ such that $\alpha(\bar{t})$ covers $S$?
Action Executability can be seen as a generalization of Plan Existence: any Plan Existence instance can be easily reduced to Action Executability, e.g., by adding a new task $t$ with new action $a$, letting $S={\{a\}}$, and adapting $\prec^{+}$ or $\delta$ to ensure that all other tasks are executed before $t$. Such direct reductions can be used to transfer algorithmic and hardness results in the respective directions; in our setting, we establish all of our hardness results for the “simpler” Plan Existence, while all of our algorithms can deal with the “harder” Action Executability.
@@
Input: A task network $tn$ in a domain $D$, an initial state $s_{0}\subseteq F$, and a target state $s_{g}\subseteq F$.
Question: Is there a solution to $(D,tn,s_{0})$ with a linearization yielding a state $s\supseteq s_{g}$?
While the considered problems are stated in their decision variants for complexity-theoretic purposes, all of our obtained algorithms are constructive and can output a solution and a corresponding linearization as a witness. Further, for primitive networks, all the problems are clearly in NP. In some results, we use the state transition graph, i.e., the multigraph whose vertices are the states reachable from the initial state $s_{0}$ by any sequence of actions, and a transition from one state to another via an action $a$ is represented by an arc labeled $a$.
3 Primitive Task Networks
Our first set of results focuses on the classical complexity of the targeted problems on primitive task networks. Before proceeding with our analysis, we make a general observation about the state space in our instances.
The State Space.
As a basic precondition for solving our problems on more sophisticated task networks, we need to ensure tractability on trivial primitive task networks where $\prec^{+}=\emptyset$ (antichains). Unfortunately, all but the first of our problems of interest are intractable even on these degenerate networks, as the hardness of Plan Existence in STRIPS planning Bylander (1994) transfers to HTN planning by an HTN-to-STRIPS translation Alford et al. (2016); see also the related work of Nebel and Bäckström Nebel and Bäckström (1994) and Erol et al. Erol et al. (1996). As discussed above, the hardness of Action Executability follows from the hardness of Plan Existence. The hardness of State Reachability follows as well: there is a trivial reduction from Plan Existence that adds a special task $t_{g}$ and adapts all actions and $s_{g}$ such that $s_{g}$ is only reachable if $t_{g}$ is executed, and $t_{g}$ is only executable after all other tasks have been executed.
A requirement for the above reductions is that the state space (i.e., the number of states in the state transition graph) can be large, and in particular not bounded by any constant. Hence, we hereinafter analyze the affected problems (State Reachability, Plan Existence, and Action Executability) in the bounded state-space setting.
The Intractability of Tree-like HTNs.
[…]
@@ 3 Primitive Task Networks
Plan Verification is NP-complete, even when restricted to primitive task networks such that $\prec^{+}$ is a collection of chains, $|A|=2$, and $\operatorname{prec}(a)=\operatorname{del}(a)=\operatorname{add}(a)=\emptyset$ for all actions $a\in A$.
@@
We reduce from Shuffle Product with the alphabet $\Sigma={\{a,b\}}$. Let $A=\Sigma$. Construct a task for each letter in the words $c_{1},\ldots,c_{w}$; let the task $t_{i,j}$ refer to the $j^{\text{th}}$ letter of the word $c_{i}$, $i\in[w]$. Let $\prec^{+}$ be such that the tasks for each word form a chain in the respective order, i.e., let $\prec$ consist of the edges $(t_{i,j},t_{i,j+1})$, for all $i\in[w]$ and $j\in[|c_{i}|-1]$. Then, $\prec^{+}$ partitions $T$ into $w$ chains (LEFT side of Fig. 3). Now, set $\alpha(t_{i,j})$ to the action representing the $j^{\text{th}}$ letter of $c_{i}$. Let the input plan $p$ be such that $p=u$. The correctness of the reduction follows since verifying whether the tasks can be arranged to match the input plan corresponds to interlacing the words to form $u$. ∎
@@
State Reachability and Plan Existence are NP-complete, even when restricted to primitive task networks such that $\prec^{+}$ is a collection of chains, $|A|\leq 5$, and the state transition graph has at most $4$ states.
Proof Sketch.
We use a similar reduction as for Theorem 1, (see Fig. 3), but encode $u$ as a separate chain and adapt the actions so that executable plans must alternate between LEFT and RIGHT tasks, and executed RIGHT tasks must be preceded by executed LEFT tasks of the same color. ∎
Figure 3: Constructed task network for reductions from Shuffle Product with words $c_{1},\ldots,c_{w}$ and $u$. Each word is represented by a chain of tasks, where each task represents a letter in that word. The two letters in the alphabet are represented in orange and blue.
Chains and Antichains.
Given the seemingly discouraging lower bounds obtained so far, one might wonder whether there are natural structural restrictions on the network which can be exploited to obtain polynomial-time algorithms. Here, we chart the route towards identifying these by first establishing the tractability of the considered problems (under the state-space restrictions justified in Sec. 3) in the extremes of $\prec^{+}$, where $T$ is a chain (total order) or antichain ($\prec^{+}\ =\emptyset$).
When $T$ is a primitive chain, it is easy to observe that all of our problems are polynomial-time solvable (see, e.g., Erol et al. (1996); Behnke et al. (2015)). Plan Verification restricted to primitive antichains is also trivial, as it suffices to check whether the plan is executable in the initial state and the multiset of actions defined by the tasks in the network matches the actions in the plan. Handling antichains for the other problems requires more effort, and no tractability results exist that only assume the state space to be bounded.
Theorem 3.
State Reachability is polynomial-time solvable on primitive antichain task networks with constantly many states in the state transition graph.
@@
Let $k$ be the number of states, $s_{0}$ the initial state, and $s_{g}$ the target state. We first compute the state transition graph ${\cal G}$ of ${\cal I}$ in quadratic time. In ${\cal G}$, if there are more than $k$ edges between any pair of states, select any $k$ of them and discard the rest. Then, consider all simple edge paths from $s_{0}$ to $s_{g}$. Note that there are $\mathcal{O}\!\left(2^{k}\cdot k^{k-1}\right)$ such paths, as there are $\mathcal{O}\!\left(2^{k}\right)$ simple vertex paths, each of length at most $k-1$, and each step might choose between up to $k$ different edges. For each simple edge path, check whether there is any action that is used more times in the path than there are tasks of that action in $T$. If there exists a simple edge path for which the number of required tasks does not exceed the number of available ones, then give a positive answer, and otherwise reject.
If such a path is found, the instance is indeed a yes-instance as there is a task sequence that results in a path to $s_{g}$ in ${\cal G}$. For the other direction, let there be a task sequence that leads to the state $s_{g}$. Consider the subsequence obtained by exhaustively removing cycles in ${\cal G}$ from that sequence. The resulting task sequence describes a simple edge path from $s_{0}$ to $s_{g}$. It might employ edges that were discarded from ${\cal G}$ by the algorithm. However, in this case there are $k$ other edges connecting the respective states in ${\cal G}$. Thus, one can choose to replace deleted edges by existing ones in such a way that each replacement action occurs at most once in the path, which ensures that there are enough tasks to actually walk this path. Hence, there is a path from $s_{0}$ to $s_{g}$ that only uses edges that are present in ${\cal G}$ after the deletion, and for which the number of required tasks does not exceed the number of available ones. As the algorithm will find this path, it will correctly give a positive output. Finally, the overall running time can be upper-bounded by $\mathcal{O}\!\left(|{\cal I}|^{2}+|A|k+|T|+2^{k}\cdot k^{k}\right)$. ∎
To obtain an analogous result for Action Executability, a more sophisticated approach is necessary, since a solution may be required to visit a state multiple times.
@@
Action Executability is polynomial-time solvable on primitive antichain task networks with constantly many states in the state transition graph.
Proof Sketch.
Action Executability on primitive antichains can be expressed by deciding whether there is a walk in the state transition graph such that each action is used in the path at least as many times as required by $S$ and at most as many times as there are respective tasks in $T$. We bound the number of actions by a function of the number $k=O(1)$ of states by identifying equivalent actions that have the same impact on the states in the state transition graph. Now we enumerate all possible types of walks in constant time by branching on a “main-path” from $s_{0}$ to any $s\supseteq s_{g}$ and which sets of simple cycles the walk uses. For all branches where the main-path and all selected cycles are connected, it remains to determine the number of times each cycle is traversed. We do so by solving an Integer Linear Program (ILP) where there are variables stating the number of traversals of each of the $O(1)$ selected cycles, and a constraint for each action requiring that the number of times it is executed (as defined by the variables and the number of times it occurs in the main-path and the respective cycles) is admissible w.r.t. $S$ and $T$. ∎
Generalized Partial Order Width.
We have seen that the considered problems are polynomial-time solvable on primitive chains and antichains. Here, we unify and extend these results by establishing tractability for primitive HTNs of constant generalized partial order width (chains and antichains have generalized partial order width $1$ and 00, respectively).
Theorem 5.
Plan Verification is polynomial-time solvable on primitive task networks of constant generalized partial order width.
Proof Sketch.
We use dynamic programming to decide, for each index $i$ in the plan and all combinations of indices $h_{1},\ldots,h_{w}$ in the $w=O(1)$ many chains partitioning the non-isolated tasks, whether there is a solution matching the first $i$ actions in the plan and using exactly the first $h_{1},\ldots,h_{w}$ tasks of each chain, respectively. Clearly, this is possible for $i=h_{1}=\ldots=h_{w}=0$ and we iteratively add a next admissable task from a chain or the isolated tasks. ∎
Theorem 6.
State Reachability and Action Executability are polynomial-time solvable on primitive task networks with constantly many states in the state transition graph and constant generalized partial order width.
Proof Sketch.
We note that the number $|E|$ of action equivalence classes that have the same impact on the state graph is constant. We use this in a dynamic programming approach similar to the one used for Theorem 5. Here, we decide for each state $s$, indices $h_{j},j\in[w]$ in the $w$ chains, and numbers of actions $r_{i},i\in[|E|]|$ in the equivalence classes, whether there is a solution with a linearization that uses exactly the first $h_{j}$ tasks of each chain $c_{j}$, results in state $s$, and uses exactly $r_{i}$ tasks of each action equivalence class $e_{i}$. We initialize all variables to be FALSE except for the one with $s=s_{0}$ and all $h_{j}$ and $r_{i}$ are 00. When a variable is set to TRUE (including the initial one), for each next index in a chain and each equivalence class for which the $h_{j}$ and $r_{i}$ imply that there is still an unused isolated task, we test whether this task can be appended to a sequence witnessing the TRUE variable. If so, the corresponding variable is set to TRUE as well. ∎
4 Hierarchical Task Networks
Considering HTNs with compound tasks adds another, seemingly opaque layer of difficulty to the considered planning problems. In this section, we show that—at least in terms of classical complexity—we can cleanly characterize the jump from primitive to compound through three measures on compound task networks: we obtain an algorithmic meta-theorem that allows to lift polynomial-time solvability from primitive to compound task networks if all three of these measures are bounded. Moreover, this result is tight in the sense that none of the three measures can be dropped. The three measures we need are ${C_{\#}}$, ${C_{s}}$, and ${C_{d}}$:
Definition 1.
For an initial task network $tn=(T,\prec^{+},\alpha)$ in a domain $D$, we refer to the number of compound tasks in $tn$ by ${C_{\#}}=|{\{t\in T\mid\alpha(t)\in{\cal C}\}}|$. The maximum size of a network that a task can be decomposed into is ${C_{s}}=\max{\{|T_{m}|\mid(c,(T_{m},\prec^{+}_{m},\alpha_{m}))\in M\}}$. The maximum depth $C_{d}$ of a network is defined recursively: ${C_{d}}=0$ for primitive task networks, and a compound task network has ${C_{d}}=i$ if $i$ is the smallest integer such that every decomposition of all compound tasks in the initial network $T$ results in a new network where ${C_{d}}\leq i-1$. We let ${C_{d}}=\infty$ if no such $i$ exists (cyclic decomposition).
Lemma 7.
If ${C_{d}}\neq\infty$, then a task network $tn=(T,\prec^{+},\alpha)$ can be decomposed into at most ${C_{c}}^{\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}}$ pairwise non-isomorphic primitive task networks $tn^{\prime}=(T^{\prime},{{\prec^{+}}}^{\prime},\alpha^{\prime})$. For each of these, $|T^{\prime}|\leq|T|+{C_{\#}}\cdot({C_{s}}^{{C_{d}}}-1)$.